OpenStudy (anonymous):

use the method of substitution to solve the system of linear equation. 6a=2-b 6a=5-b

OpenStudy (jiteshmeghwal9):

This is wrong system of linear equations in two variables since 6a should be equal to any common equation. If 6a=2-b then 6a can't be equal to 5-b & if \(6a=5-b\) then 6a can't be equal to 2-b.

OpenStudy (anonymous):

so there is no solution?

OpenStudy (jiteshmeghwal9):

yes :)

OpenStudy (anonymous):

ok